You own a portfolio that has $3.800 invested in stocks and $5,700 invested in bonds. What is the expected return of the porfolio if stocks and bonds are expected to yield a return of 11% and 9%, respectively? (Round your answer to 2 decimal places.) Expected return
Answer #1

weight of stock A w1= stockA/(stoack A + stock B) = 3800/(3800+5700) = 0.4

weight of stock B w2= stockB/(stoack A + stock B) = 5700/(3800+5700) = 0.6

expected returns:

r1 = 11%

r2 = 9%

Expected return on portfolio = w1*r1 + w2*r2 = 0.4*11 + 0.6*9 = 9.8%